# DDIL: Improved Diffusion Distillation with Imitating Learning

### v1

@@ C.1 Progressive Distillation
Progressive Step distillation aims at reducing the number of timesteps $T$ of the sampling (reverse) process in the diffusion models by learning a new student model. Starting from timestep $t$ within the reverse diffusion process, given the discretization interval $k$, $N$ steps of the teacher are distilled into $N/k$ steps of the student ($T=N$ in the first iteration of step-distillation) Salimans & Ho (2021). We query the teacher model at timesteps $t-k$ and $t-2k$ while the student estimates are obtained at $t-2k$ using a DDIM Song et al. (2020).
Following the formulation in Li et al. (2023b), the teacher model is unrolled for two DDIM steps to $t-k$ and $t-2k$ starting at timestep $t\in\left[T\right]$ and $0\leq t-2k<{t-k}$ with input noisy latent $\mathbf{z}_{t}$ while student model performs one denoising step. Where $\hat{\mathbf{v}}_{t}^{\text{s}}$ is the velocity estimate from the student model $\hat{\mathbf{v}_{\eta}(\mathbf{z}_{t},t)}$. The student model predicts the latent $\mathbf{z}_{t-2k}^{\text{s}}$ from $\mathbf{z}_{t}$ of the teacher and thus $\mathbf{z}_{t-2k}^{\text{s}}=\mathbf{z}_{t-2k}$. Progressive distillation loss with guidance conditioned teacher is denoted by,
@@
Here, the student model is trained with teacher forcing as is evident in Equation 3. During sampling from the student model, the teacher observations are not provided and therefore, the student model can drift from the expected trajectory Huszár (2015).
@@ Appendix E Co-variate shift Analysis
To validate our hypothesis of covariate shift from accumulation of error, we conduct a mixed-rollout evaluation using a 32-step CFG teacher and a 4-step DDIL-distilled student on the MS-COCO 2017 (5K) dataset. Both models achieve similar FID scores ( 22.5) with the teacher model having a CLIP score of 0.321. We assume a 32-step teacher model, where 1 student step is equivalent to 8 teacher steps. This allows for alignment between teacher and student estimates at specific timesteps in the diffusion process ${999,749,500,250}$. This setting enables stochastic mixing between the teacher and student models during inference by choosing a state-less DDIM solver. We investigate three settings to assess if the teacher model can improve student generation from intermediate time steps. We vary the prior probability $(p_{T})$ of selecting the teacher model for each (t-250) of transition. Results in Tab. 7 show that decreasing $p_{T}$ (less frequent teacher usage) leads to a decline in CLIP score, suggesting the teacher model improves student predictions across various time steps. To understand if there are any critical timesteps for teacher intervention, we analyze the impact of switching between teacher and student at specific time steps (Tab. 7 and 7). The results indicate that teacher intervention, either early or late in the diffusion process, can improve generation quality compared to student-only inference. This further supports the presence of covariate shift and its impact on student model.
@@ E.1 Covariate Shift Visualization
Appendix F Prompt Inversion
In Table 8, we investigate if the underlying map from ‘noise’ space to ‘data’ space is preserved during the distillation phase of diffusion. Updating the map could have an implication on adopting various tools obtained on pre-trained diffusion model with applications in personalization etc. To capture this we consider a setting where we obtain inverted prompts of a reference image using Mahajan et al. (2023) on the COCO dataset and then pass the inverted prompts to distilled models to capture the similarity of the generated image to the reference image and also diversity of generations. As PH2P returns the optimal token for a given image, if a relative change in the map is minimal we expect the generated output to be more aligned with the reference image. If distilled model has good behavior we expect distilled models to preserve diversity of generation on invertd prompts too, our overall findings are consistent with text-guided generation for inverted prompts too.

### v2

@@ C.1 Progressive Distillation
Progressive Step distillation aims at reducing the number of timesteps $T$ of the sampling (reverse) process in the diffusion models by learning a new student model. Starting from timestep $t$ within the reverse diffusion process, given the discretization interval $k$, $N$ steps of the teacher are distilled into $N/k$ steps of the student ($T=N$ in the first iteration of step-distillation) [32]. We query the teacher model at timesteps $t-k$ and $t-2k$ while the student estimates are obtained at $t-2k$ using a DDIM [36].
Following the formulation in [15], the teacher model is unrolled for two DDIM steps to $t-k$ and $t-2k$ starting at timestep $t\in\left[T\right]$ and $0\leq t-2k<{t-k}$ with input noisy latent $\mathbf{z}_{t}$ while student model performs one denoising step. Where $\hat{\mathbf{v}}_{t}^{\text{s}}$ is the velocity estimate from the student model $\hat{\mathbf{v}_{\eta}(\mathbf{z}_{t},t)}$. The student model predicts the latent $\mathbf{z}_{t-2k}^{\text{s}}$ from $\mathbf{z}_{t}$ of the teacher and thus $\mathbf{z}_{t-2k}^{\text{s}}=\mathbf{z}_{t-2k}$. Progressive distillation loss with guidance conditioned teacher is denoted by,
@@
Here, the student model is trained with teacher forcing as is evident in Equation 2. During sampling from the student model, the teacher observations are not provided and therefore, the student model can drift from the expected trajectory [8].
@@ Appendix E Co-variate shift Analysis
To validate our hypothesis of covariate shift from accumulation of error, we conduct a mixed-rollout evaluation using a 32-step CFG teacher and a 4-step DDIL-distilled student on the MS-COCO 2017 (5K) dataset. Both models achieve similar FID scores ( 22.5) with the teacher model having a CLIP score of 0.321. We assume a 32-step teacher model, where 1 student step is equivalent to 8 teacher steps. This allows for alignment between teacher and student estimates at specific timesteps in the diffusion process ${999,749,500,250}$. This setting enables stochastic mixing between the teacher and student models during inference by choosing a state-less DDIM solver. We investigate three settings to assess if the teacher model can improve student generation from intermediate time steps. We vary the prior probability $(p_{T})$ of selecting the teacher model for each (t-250) of transition. Results in Tab. 9 show that decreasing $p_{T}$ (less frequent teacher usage) leads to a decline in CLIP score, suggesting the teacher model improves student predictions across various time steps. To understand if there are any critical timesteps for teacher intervention, we analyze the impact of switching between teacher and student at specific time steps (Tab. 9 and 9). The results indicate that teacher intervention, either early or late in the diffusion process, can improve generation quality compared to student-only inference. This further supports the presence of covariate shift and its impact on student model.
@@ E.1 Covariate Shift Visualization
From Fig. 5, we can observe that more we use teacher model structural coherence is improved of the horse. More specifically if we switch from teacher to student from 750 i.e., run teacher model to provide initialization then run student model the structure is largely preserved. So this experiment shows importance of early denoising steps compared to later iterations within denoising process.
Alternatively if we switch from student to teacher i.e., see if the teacher has ability to correct and generate good image, we can observe from Fig 5 if we switch to teacher after t=500 i.e, half of denoising steps teacher cannot fix already in-coherent structure of horse’s legs (back). This experiment also illustrates potential for non-uniform denoising for improved quality vs latency tradeoff in diffusion models.
@@
Appendix F Prompt Inversion
In Table 10, we investigate if the underlying map from ‘noise’ space to ‘data’ space is preserved during the distillation phase of diffusion. Updating the map could have an implication on adopting various tools obtained on pre-trained diffusion model with applications in personalization etc. To capture this we consider a setting where we obtain inverted prompts of a reference image using [24] on the COCO dataset and then pass the inverted prompts to distilled models to capture the similarity of the generated image to the reference image and also diversity of generations. As PH2P returns the optimal token for a given image, if a relative change in the map is minimal we expect the generated output to be more aligned with the reference image. If distilled model has good behavior we expect distilled models to preserve diversity of generation on invertd prompts too, our overall findings are consistent with text-guided generation for inverted prompts too.
@@ Appendix G Additional Qualitative Examples
In this section, we compare various trajectory based distillation techniques on SDv1.5 and show efficacy of DDIL(+PD) in terms of quality, diversity compared to other publicly available distilled checkpoints from Instaflow, ADD (SDv2.1) and LCM compared to pre-trained diffusion model SDv1.5.